OpenStudy (anonymous):

Lonnie places $950 in a savings account that earns 5.75% simple interest. Find the total amount in the account after four years.

OpenStudy (anonymous):

total amount = P+SI =950 + PTR/100 =950 + 950*4*5.75/100 =1168.5

OpenStudy (anonymous):

you wont get richer with this rate of simple interest

OpenStudy (anonymous):

William’s inheritance from his great uncle came to $225,000 after taxes. If William invests this money in a savings account at 7.3% interest, how much will he earn from the account each year?